1)What is the pH of a solution in which 296 mL of HCl(g), measured at 26.4 ∘C and 1.02 atm, is dissolved in 1.5 L of aqueous solution?

2)What volume of a concentrated HCl solution, which is 36.0% HCl by mass and has a density of 1.179 g/mL, should be used to make 5.30 L of an HCl solution with a pH of 1.6?

Determine the pH of an HF solution of each of the following concentrations. In which cases can you not make the simplifying assumption that x is small? (Ka for HF is 6.8×10−4.)

1) 0.300 M
2) 5.10×10−2 M
3) 2.50×10−2 M
1
Expert's answer
2020-04-03T11:48:52-0400

a) pV=nRT n= pV/RT = 1.02 x 0.296 / 0.082 x 299.4 = 0.0123 mol and 0.0123/1.5 =0.00819851087 M.

pH= 2.086.

b) antilog(-1.6)= 0.0251 M.

5.3 x 0.0251=0.13303moles.

0.13303 x 36.5 =4.855595 g

4.855595x100/ 36.0 = 13.4877639 g.

13.4877639/1.179 = 11.4400033 ml.

1) 0.522878745

2) 1.3

3) 1.6